OpenStudy (anonymous):

if a spherical tank 2 m in diameter can be filled with a liquid for $300, find the cost to fill the tanks of 6 diameter.

OpenStudy (anonymous):

V of small sphere = 4/3*pi*(d/2)^3=4/3*pi*d^3/8 V of big sphere = 4/3*pi*(3d/2)^3=4/3*pi*27d^3/8 thus the volume of the big sphere is 27 times the volume of the small sphere the cost for filling the tank of 6 diameter = 27*300=$8100